If it takes a Dell Dimension-8300 computer 4 hours to crack a UNIX password file and Gateway E-6100 computer
twice as long to crack the same password file, how long would it take the two of them, working together
(assume we can get two of them to work together) to crack the file, if each computer works at its respective
individual speed?

Answer by Earlsdon(6294)

If the Dell computer takes 4 hrs to crack the password file, then it can crack 1/4 of the file in 1 hour.
If the Gateway computer takes 2(4) = 8 hrs to crack the same password file, then it can crack 1/8 of the file in 1 hour.
Working together, the two computers can crack (1/4 + 1/8 = 3/8) of the file in 1 hour.
Therefore, it would take them 8/3 hours, working together to crack the entire password file.
8/3 hours = 2 hours 40 minutes.